Operation Management

Brandon Production is a small firm focused on the assembly and sale of custom  computers. The firm is facing stiff competition from low-priced alternatives, and is  looking at various solutions to remain competitive and profitable. Current financials  for the firm are shown in the table below. In the first option, marketing will increase  sales by 50%. The next option is vendor (supplier) changes, which would result in  a decrease of 10% in the cost of inputs. Finally there is an Operation Management option, which would reduce production costs by 25%. Which of the options would  you recommend to the firm if it can only pursue one option? In addition, comment  on the feasibility of each option. Business Function Current Value Cost of Inputs $50,000 Production Costs $25,000 Revenue $80,000 [7 Marks]
Answer #1

ANSWER :


Option 1 :


Increase of sales by 50% will result into :


Revenue increase = 80000*0.5 = 40000 ($)


Profit increase = 40000 - (50000 + 25000)/2 

= 2500 ($)


50% increase in sales by marketing may be difficult but is feasible. 


Option 2 :


increase in profit  = Decrease in input costs

= 50000*0.10

= 5000 ($) 


By vendor management, it is feasible.


Option 3 :


Profit increase = Decrease in production costs

= 25000 * 0.15

= 3750 ($) 


By proper operation management it is feasible.


From the above, we note that maximum benefit can be achieved by vendor management and decreasing input costs by 10%. .



answered by: Tulsiram Garg